OpenStudy (anonymous):

simplify 14-7x over x-2

OpenStudy (anonymous):

isnt it -7?

OpenStudy (anonymous):

\[\frac{14-7x}{x-2}=\frac{7(2-x)}{x-2}=7\times \frac{2-x}{x-2}=7\times (-1)\]

OpenStudy (anonymous):

yes, it is \(-7\)

OpenStudy (anonymous):

Thank you for helping me!!

OpenStudy (anonymous):

yw